# Modelling Gaia photometry signals of dark halos

### v1

@@ Abstract
We use the framework of microlensing to show that observations of binary systems, such as those made by Gaia, combined with follow-up weak lensing measurements, can provide a means to probe halos of exotic matter, possibly clumped around compact objects such as black holes. This could potentially cover a broad range of physical scenarios - from dark matter mini-halos to black holes with bosonic configurations around them, known as hair. Assuming that light can freely propagate through the halo of the exotic matter, the companion star will produce characteristic, sizable lensing signatures due to the deviation from a central gravitational potential. The signature of the multiple images, the magnification and the light-curve could be the smoking gun of such structures. We discuss how the precise observations of the Gaia survey offer an opportunity to search for new, yet undiscovered fundamental fields interacting gravitationally with baryons.
keywords:
keyword1 – keyword2 – keyword3
pubyear: 2025pagerange: Modelling Gaia photometry signals of dark halosA
1 Introduction
@@
Indeed, a new pressure-less (dust-like) component of matter in the Standard Model of particle physics is required by a plethora of observational probes at larger or smaller scales in the Universe, e.g. the Cosmic Microwave Background, galactic rotation curves, or supernovae observations (see e.g. Famaey & McGaugh, 2012; Hui et al., 2017; Roszkowski et al., 2018). At the same time, theories beyond the Standard Model typically predict the existence of new particles of various masses and spins. Within an astrophysical context, this begs the question of whether stable configurations of such exotic matter can form, be it dark matter or other fundamental fields. Clumps of exotic matter could form either isolated or around compact objects such as black holes (BHs), through either the process of gravitational collapse (Gondolo & Silk, 1999; Schunck & Mielke, 2003; Sadeghian et al., 2013; Brito et al., 2016; Pombo & Saltas, 2023), or the so–called mechanism of scalarisation/vectorisation (Cardoso et al., 2013; Silva et al., 2018; Dima et al., 2020; Cunha et al., 2019; Dolan, 2018; Pani et al., 2012). The latter intimately relates to the question of the nature of BHs, and the existence of hair-like configurations around them (Liebling & Palenzuela, 2023; Herdeiro et al., 2021; Berti et al., 2021). Our proposed probe will set out a promising and powerful tool to detect exotic structures in our galaxy, as well as place constraints on the parameters of their profiles from observations.
In this work, we introduce a framework to probe the existence of stable configurations of clumps of exotic matter based on the combination of astrometry and lensing observations of binary systems within our galaxy. The astrometry frontier is currently led by the high-precision observations of Gaia (Gaia Collaboration et al., 2016, 2023a, 2023b), and promises to drastically improve in the future. On the other hand, several precision surveys aiming at (micro-)lensing are underway while others have already been completed (Alcock et al., 2000; Tisserand et al., 2007; Udalski et al., 2015).
@@
Under certain conditions, particles beyond the Standard Model of any spin, which may or may not contribute to dark matter, can form halos at astrophysical scales. This may be the result of collapse under their own gravity, or through clumping around some compact object such as a BH  [^2] The evolutionary channels forming such systems play a key role but will not be the topic of this work.. This includes the case of the formation of dark matter structures at small (sub-galactic) scales, the bosonic condensates known as boson stars, or the fundamental question of BH physics, that is, whether BHs may develop “hair" configurations around them made out of some exotic field.
@@
The key novelty of our work concerns on applying the microlensing framework on searching for dark haloes (DHs) with Gaia-like binaries (involving the relevant masses, binary sizes, etc.). Our discussion will cover all dark matter theories or scenarios where exotic matter is predicted to form clumps of matter at astrophysical scales, irrespective of the formation channel. The fine details of the profile characterising the scaling of the density of the halo with the radius is certainly a model-dependent quantity. In practice, a parametrization is enough to capture all the essential features exhibited in full-fledged numerical simulations, a notable example being the Navarro-Frenk-White profile for dark matter (Navarro et al., 1996). Our proposed probe will set out a promising tool to place constraints on such profiles from small-scale astrophysical observations. It will complement and extend similar previous works where the lensing signatures of (isolated) exotic matter configurations were studied assuming the thin-lens approximation (Croon et al., 2020a, b; Croon & Munoz, 2024; Romao & Croon, 2024). One of the key aspects of our work is the focus on binary systems which are within the reach of the Gaia survey. Such binaries can be observed photometrically by both Gaia as well as ground-based photometry telescopes. The combination of high-precision astrometry with photometric observations offers a highly promising framework for the detection and analysis of such systems.
@@
As we will see, for binary systems such as those recently discovered by Gaia (BH1 El-Badry et al. (2022), BH2 El-Badry et al. (2023), BH3 Gaia Collaboration (2024)) the existence of the DH can leave distinct signatures in the lensing signal, within the precision of current instruments. This is predominantly imprinted on the structure of the multiple images and the resulting magnification or the light curve, respectively. Our analysis shows that the change in apparent magnitude of the star during a transit (i.e., when the star enters the “shadow” of the DH lens), depending on the geometry, can be as large as a few magnitudes during ingress/egress or a few tenths of a magnitude during the rest of the transit. The proposed framework relies on the combination of Gaia’s astrometric measurements with follow-up lensing observations of the binary system. This synergy is key - the precise reconstruction of the orbital parameters from Gaia will not only allow an accurate follow up of the binary, but will also break degeneracies with lensing observables.
@@
We structure the paper as follows: In Section 2 we introduce the theoretical framework for weak lensing by a DH, and in Sections 3, 4, and 5, we discuss the key phenomenological imprints and population estimates of such systems. We summarise in Section 6.
2 Weak lensing through a dark halo
[…]
@@ 2 Weak lensing through a dark halo
Figure 1: A schematic representation of our setup. The gravitational lens (DH) is a self-gravitating clump of exotic matter with or without a BH at the centre. The light source is a stellar object such as Gaia’s BH1 binary system. Propagation of light through the density of the DH induces non-trivial signatures of multiple images and magnification, which can be observed through photometric observations. Gaia’s astrometry and photometry observations allow us to identify such binary systems in our galaxy and probe a broad range of theories for dark matter and other particles beyond the Standard Model. The impact parameter ($b$) and deflection angle ($\alpha$) are explained in the text.
@@
where the integration is taking place along the path of the light-ray ($x$). Here, $b$ is the impact parameter of the ray (see also Fig. 1) and $M(x;b)$ is the effective mass profile, while $r=\sqrt{x^{2}+b^{2}}$. When $b\leq R$, $M(x;b)$ is given by $M(x;b)=\frac{GM}{2R^{3}}\left(b^{2}+x^{2}\right)^{3/2}$, while for $b>R$ it is a constant, $M(x;b)=-GM$. We can then express the deflection angle $\alpha$ as
$\alpha=\left\{\begin{tabular}[]{cc}$\frac{4GM}{bc^{2}}$&$b>R$\\ $\frac{4GM}{bc^{2}}\left(1-\frac{a^{3}}{R^{3}}\right)$&$b\leq R$\end{tabular},\right.$ (4)
where $a^{2}=R^{2}-b^{2}$. This has also been derived by Croon et al. (2020a) (see also Croon et al. (2020a, b); Croon & Munoz (2024); Romao & Croon (2024)). We can define the characteristic quantity, $\alpha_{o}\equiv\frac{GM}{c^{2}R}\approx 0.425^{\prime\prime}\times\left(\frac{R_{\odot}}{R}\right)\left(\frac{M}{M_{\odot}}\right),$ where for a lens with the mass and radius of the Sun, we recover a deflection angle $\alpha_{\odot}=4GM_{\odot}/c^{2}R_{\odot}\approx 1.7^{\prime\prime}$. Since neither the light source nor the observer is at infinity with respect to the lens, a correction factor needs to be applied, which has the form
@@
where $D_{\rm LS}$ is the lens-source, and $D_{\rm S}$ the observer-source distance respectively. We will further denote the observer-lens distance with $D_{L}$. The ratio $\frac{D_{\rm LS}}{D_{\rm S}}$ is essentially the ratio of the size of the binary system over the distance to the observer. To get an idea, if the distance of the star from the lens is $\sim 1.4AU\approx 300R_{\odot}$, while the distance of the system from the observer is $D_{\rm S}=480pc\approx 2.127\times 10^{10}R_{\odot}$, this implies that $D_{\rm LS}/D_{\rm S}\approx 1.4\times 10^{-8}$.
3 Dark halo light curve (no central black hole)
@@
$u=x-\frac{1}{x}\left\{\begin{tabular}[]{cc}$1$&$,x\geq x_{R}$\\ $1-\left[1-(x/x_{R})^{2}\right]^{3/2}$&$,x<x_{R}$\end{tabular}.\right.$ (7)
@@
where we have redefined $\tilde{\alpha}=\bar{\alpha}/\theta_{E}$. This essentially corresponds to looking for the $x$-value of the intersections of the lines $y=x-u$ (l.h.s) with the curve $y=\tilde{\alpha}(x)$ (r.h.s) for different values of the normalised true position of the luminous source (i.e., $u$).
@@
This process is shown in Figure 2 for DHs of various sizes. As can be seen, for every value of $u$ there exist either one or three intersections between the lines of $y=x-u$ and the respective curves of $y=\tilde{\alpha}(x)$, which correspond to one or three images of the light source created by the DH lens. It is worth noticing that varying $u$ corresponds to evolving the binary system in time. One can show that the slope of $\tilde{\alpha}$ at $x=0$ is $\left.\frac{d\tilde{\alpha}}{dx}\right|_{x=0}=\frac{3}{2x_{R}^{2}}$, which means that when the size of the lens is $x_{R}\geq\sqrt{3/2}$ the slope is less than $1$ and there can be a maximum of one image since there can be only one intersection between $y=x-u$ and $y=\tilde{\alpha}(x)$, while when $x_{R}<\sqrt{3/2}$ there can be as many as $3$ (see also Croon et al., 2020a, for discussion on this). In general, when the true position of the light source is far from the axis connecting the observer to the lens (i.e., large values of $u$) there is only one image produced by the DH lens from the intersection of $y=x-u$ with $y=1/x$. We remind here that, in this case, we consider a DH without a point-like mass (BH) at its centre. We will get back to this issue later.
@@
Figure 3: The change in magnitude $\Delta m$ of a light source during an eclipse by a DH lens of size $x_{R}=0.7$, as a function of the true position of the source on the sky, the latter parametrised by the angle $\beta$ (see equation (12)). We assume that the lensing potential is solely produced by the DH, i.e., no BH exists at the centre of the DH. We remind that $\Delta m$ is directly related to the total magnification $\mu$, defined in equation (12). Our results appear to be in qualitative agreement with Romao & Croon (2024).
@@
where $v_{\perp}$ corresponds to the relative velocity of the source-lens in the direction perpendicular to the line-of-sight of the observer. The presence of a DH would affect $\theta_{E}$, and possibly the velocity $v_{\perp}$ if the trajectory of the source is modified due to the difference in the gravitational potential and/or the existence of dynamical friction effects. We will not explore here such effects, and leave them for future research. It is important to notice that the physical quantities which enter into the Einstein time cannot be measured by lensing alone, and more information is needed. Astrometry observations can indeed break the degeneracy between these parameters through a measurement of the velocity. To model the actual light curve associated with the lensing effect in the time-domain we assume a trajectory for the source as
@@
Assuming a Keplerian orbit for the light source around the DH lens (i.e., $v_{\perp}\approx\sqrt{\frac{GM}{A}}$, with $A$ the semi-major axis of the orbit which is $A\approx D_{LS}$), the Einstein time is $t_{E}=2D_{LS}/c$. Using Kepler’s laws the Einstein time can be expressed as a function of the period $T$ and the total mass of the system $M$ as
@@
For a system like the Gaia BH1, the total mass is $\sim 11M_{\odot}$, the mass ratio between the total mass and the luminous-star mass is $\sim 14$, the period is $185.6$ d, and the size of the orbit is $\sim 1.4AU$ (El-Badry et al., 2022). Then, it turns out that the Einstein time is of the order of $\sim 23$ min. The time evolution of the apparent magnitude of the source as it passes behind the lens can be seen in Figure 4 for the case of two occultation events, the first with $u_{0}=0$ and the second with $u_{0}=0.91x_{R}$, respectively.
Figure 4: Similar $\Delta m$ to Figure 3, but now as a function of time. The evolution depends on the characteristic time $t_{E}$ (see (13)), and we have assumed the same characteristics as those of the Gaia BH1 system. The figure shows two transit events, one with initial impact parameters $u_{0}=0$ and $u_{0}=0.91x_{R}$, respectively. The transit for the first event lasts a total of $\sim 40$min, while the spikes in the magnification that happen as the source passes behind the edge of the DH lens last $\sim 5$min each. For the rest of the transit, the source is brighter than its normal brightness and behaves like a point mass. The second transit lasts less time, $\sim 25$min since the light source grazes the lens. In this case, we observe two magnification spikes as the source goes behind the lens, while the source remains magnified throughout the transit. This plot should be compared with similar ones for e.g. by Romao & Croon (2024).
4 Dark halo light curve (Black hole with a dark halo)

### v2

@@ Abstract
We use the framework of microlensing to show that observations of binary systems, such as those made by Gaia, combined with follow-up weak lensing measurements, can provide a means to probe halos of exotic matter, possibly clumped around compact objects such as black holes. This could potentially cover a broad range of physical scenarios - from dark matter mini-halos to black holes with bosonic configurations around them, known as hair. In a dark halo-regular star binary system, assuming that light can freely propagate through the halo of the exotic matter, the companion star will produce characteristic, sizable lensing signatures due to the lens’s potential deviating from a point mass gravitational potential. The signature of the multiple images, the magnification and the light-curve could be the smoking gun of such a dark halo structure. We discuss how the precise observations of the Gaia survey offer an opportunity to search for new, yet undiscovered fundamental fields interacting gravitationally with baryons.
I Introduction
@@
Indeed, a new pressure-less (dust-like) component of matter in the Standard Model of particle physics is required by a plethora of observational probes at larger or smaller scales in the Universe, e.g. the Cosmic Microwave Background, galactic rotation curves, or supernovae observations (see e.g. Famaey and McGaugh, 2012; Hui et al., 2017; Roszkowski et al., 2018). At the same time, theories beyond the Standard Model typically predict the existence of new particles of various masses and spins. Within an astrophysical context, this begs the question of whether stable configurations of such exotic matter can form, be it dark matter or other fundamental fields. Clumps of exotic matter could form either isolated or around compact objects such as black holes (BHs), through either the process of gravitational collapse (Gondolo and Silk, 1999; Schunck and Mielke, 2003; Sadeghian et al., 2013; Brito et al., 2016; Pombo and Saltas, 2023), or the so–called mechanism of scalarisation/vectorisation (Cardoso et al., 2013; Silva et al., 2018; Dima et al., 2020; Cunha et al., 2019; Dolan, 2018; Pani et al., 2012). The latter intimately relates to the question of the nature of BHs, and the existence of hair-like configurations around them (Liebling and Palenzuela, 2023; Herdeiro et al., 2021; Berti et al., 2021). Our proposed probe will set out a promising and powerful tool to detect exotic structures in our galaxy, as well as place constraints on the parameters of their profiles from observations.
In this work, we introduce a framework to probe the existence of stable configurations of clumps of exotic matter based on the combination of astrometry and lensing photometry observations of binary systems within our galaxy, where the binary system is comprised of such a dark exotic matter configuration and a regular star that is lensed by its dark companion. The astrometry frontier is currently led by the high-precision observations of Gaia (Gaia Collaboration et al., 2016; Gaia Collaboration et al., 2023a, b), and promises to drastically improve in the future. Furthermore, Gaia (spectro-)photometry also provides a unique all sky multi-epoch survey. On the other hand, several precision surveys aiming at (micro-)lensing are underway while others have already been completed (Alcock et al., 2000; Tisserand et al., 2007; Udalski et al., 2015). Photometric self-lensing in binary systems was first predicted by Maeder Maeder (1973) and observed four decades later by Kruse and Agol Kruse and Agol (2014). The periodic nature of the phenomenon not only facilitates its detection but also opens the possibility for dedicated follow-up observations, enabling detailed characterisation of the system’s orbital and physical parameters, potentially including the halo properties of the black hole or degenerate star.
@@
Under certain conditions, particles beyond the Standard Model of any spin, which may or may not contribute to dark matter, can form halos at astrophysical scales. This may be the result of collapse under their own gravity, or through clumping around some compact object such as a BH  [^2] The evolutionary channels forming such systems play a key role but will not be the topic of this work.. This includes the case of the formation of dark matter structures at small (sub-galactic) scales, the bosonic condensates known as boson stars, or the fundamental question of BH physics, that is, whether BHs may develop “hair” configurations around them made out of some exotic field.
@@
The key novelty of our work concerns on applying the microlensing framework on searching for dark haloes (DHs) with Gaia-like binaries (involving the relevant masses, binary sizes, etc.). Our discussion will cover all dark matter theories or scenarios where exotic matter is predicted to form clumps of matter at astrophysical scales, irrespective of the formation channel. The fine details of the profile characterising the scaling of the density of the halo with the radius is certainly a model-dependent quantity. In practice, a parametrization is enough to capture all the essential features exhibited in full-fledged numerical simulations, a notable example being the Navarro-Frenk-White profile for dark matter (Navarro et al., 1996). Our proposed probe will set out a promising tool to place constraints on such profiles from small-scale astrophysical observations. It will complement and extend similar previous works where the lensing signatures of (isolated) exotic matter configurations were studied assuming the thin-lens approximation (Croon et al., 2020a, b; Croon and Munoz, 2024; Romao and Croon, 2024). One of the key aspects of our work is the focus on binary systems which are within the reach of the Gaia survey. Such binaries can be observed photometrically by both Gaia as well as ground-based photometry telescopes doing follow-up observations. The combination of high-precision astrometry with photometric observations offers a highly promising framework for the detection and analysis of such systems.
@@
As we will see, for binary systems such as those recently discovered by Gaia (BH1 El-Badry et al. (2022); Chakrabarti et al. (2023), BH2 El-Badry et al. (2023), BH3 Gaia Collaboration (2024)) the existence of the DH can leave distinct signatures in the lensing signal, within the precision of current instruments. This is predominantly imprinted on the structure of the multiple images and the resulting magnification or the light curve, respectively. Our analysis shows that the change in apparent magnitude of the star during a transit (i.e., when the star enters the “shadow” of its DH lens companion), depending on the geometry, can be as large as a few magnitudes during ingress/egress or a few tenths of a magnitude during the rest of the transit. The proposed framework relies on the combination of Gaia’s astrometric measurements with follow-up lensing observations of the binary system. This synergy is key - the precise reconstruction of the orbital parameters from Gaia will not only allow an accurate follow up of the binary, but will also break degeneracies with lensing observables.
@@
We structure the paper as follows: In Section II we introduce the theoretical framework for weak lensing by a DH, and in Sections III, IV, and V, we discuss the key phenomenological imprints and population estimates of such systems. We summarise in Section VI.
II Weak lensing through a dark halo
Here we will introduce the main equations we will use, summarising previous results in the literature (Congdon and Keeton, 2018). In the weak-field approximation, gravitational lensing can be described in terms of Fermat’s principle, where spacetime has an effective index of refraction, $n=\frac{c}{c^{\prime}}=1-\frac{2\Phi}{c^{2}},$ where $\Phi$ satisfies the Poisson equation $\nabla^{2}\Phi=4\pi G\rho$. By extremising the length of the path of a light-ray, $\delta\int_{A}^{B}\frac{c}{n}dt=0$, one arrives at an expression for the deflection angle, which is given as
$\vec{\alpha}=-\frac{2}{c^{2}}\int\vec{\nabla}_{\perp}\Phi d\lambda,$ (1)
@@ II Weak lensing through a dark halo
Figure 1: A schematic representation of our binary setup. The gravitational lens (DH) is a self-gravitating clump of exotic matter with or without a BH at the centre. The light source in orbit is a stellar object such as the one in Gaia’s BH1 binary system. Propagation of light through the density of the DH induces non-trivial signatures of multiple images and magnification, which can be observed through photometric observations. Gaia’s astrometry and photometry observations allow us to identify such binary systems in our galaxy and probe a broad range of theories for dark matter and other particles beyond the Standard Model. The impact parameter ($b$) and deflection angle ($\alpha$) are explained in the text.
@@
$\alpha=\frac{2}{c^{2}}\int_{-\infty}^{+\infty}\frac{\partial}{\partial b}\left(\frac{GM(x;b)}{\sqrt{b^{2}+x^{2}}}\right)dx,$ (3)
where the integration is taking place along the path of the light-ray ($x$). Here, $b$ is the impact parameter of the ray (see also Fig. 1) and $M(x;b)$ is the effective mass profile, while $r=\sqrt{x^{2}+b^{2}}$. When $b\leq R$, $M(x;b)$ is given by $M(x;b)=\frac{GM}{2R^{3}}\left(b^{2}+x^{2}\right)^{3/2}$, while for $b>R$ it is a constant, M(x;b)=GMM(x;b)=-GM. We can then express the deflection angle $\alpha$ as
@@
where $a^{2}=R^{2}-b^{2}$. This has also been derived by Croon et al. (2020a) (see also Croon et al. (2020a, b); Croon and Munoz (2024); Romao and Croon (2024)). We can define the characteristic quantity, $\alpha_{o}\equiv\frac{GM}{c^{2}R}\approx 0.425^{\prime\prime}\times\left(\frac{R_{\odot}}{R}\right)\left(\frac{M}{M_{\odot}}\right),$ where for a lens with the mass and radius of the Sun, we recover a deflection angle $\alpha_{\odot}=4GM_{\odot}/c^{2}R_{\odot}\approx 1.7^{\prime\prime}$. Since neither the light source nor the observer is at infinity with respect to the lens, a correction factor needs to be applied, which has the form
@@
where $D_{\rm LS}$ is the lens-source, and $D_{\rm S}$ the observer-source distance respectively. We will further denote the observer-lens distance with $D_{L}$. The ratio $\frac{D_{\rm LS}}{D_{\rm S}}$ is essentially the ratio of the size of the binary system over the distance to the observer. To get an idea, if the distance of the star from the lens is $D_{\rm LS}\sim 1.4AU\approx 300R_{\odot}$, while the distance of the system from the observer is $D_{\rm S}=480pc\approx 2.127\times 10^{10}R_{\odot}\approx D_{\rm L}$, this implies that $D_{\rm LS}/D_{\rm S}\approx 1.4\times 10^{-8}$. The numbers here are inspired by the BH1 system (El-Badry et al., 2022) and in our discussion we will often use the parameters of this system as a template for estimates.
III Dark halo light curve (no central black hole)
@@
[…]
@@
where we have redefined $\tilde{\alpha}=\bar{\alpha}/\theta_{E}$. This essentially corresponds to looking for the $x$-value of the intersections of the lines y=xuy=x-u (l.h.s) with the curve $y=\tilde{\alpha}(x)$ (r.h.s) for different values of the normalised true position of the luminous source (i.e., $u$).
@@
This process is shown in Figure 2 for DHs of various sizes. As can be seen, for every value of $u$ there exist either one or three intersections between the lines of y=xuy=x-u and the respective curves of $y=\tilde{\alpha}(x)$, which correspond to one or three images of the light source created by the DH lens. It is worth noticing that varying $u$ corresponds to evolving the binary system in time. One can show that the slope of $\tilde{\alpha}$ at $x=0$ is $\left.\frac{d\tilde{\alpha}}{dx}\right|_{x=0}=\frac{3}{2x_{R}^{2}}$, which means that when the size of the lens is $x_{R}\geq\sqrt{3/2}$ the slope is less than $1$ and there can be a maximum of one image since there can be only one intersection between y=xuy=x-u and $y=\tilde{\alpha}(x)$, while when $x_{R}<\sqrt{3/2}$ there can be as many as $3$ (see also Croon et al., 2020a, for discussion on this). In general, when the true position of the light source is far from the axis connecting the observer to the lens (i.e., large values of $u$) there is only one image produced by the DH lens from the intersection of y=xuy=x-u with $y=1/x$. We remind here that, in this case, we consider a DH without a point-like mass (BH) at its centre. We will get back to this issue later.
@@
Figure 3: The change in magnitude $\Delta m$ of a light source during an eclipse by a DH lens of size $x_{R}=0.7$, as a function of the true position of the source on the sky, the latter parametrised by the angle $\beta$ (see equation (12)). We assume that the lensing potential is solely produced by the DH, i.e., no BH exists at the centre of the DH. We remind that $\Delta m$ is directly related to the total magnification $\mu$, defined in equation (12). Our results appear to be in qualitative agreement with Romao and Croon (2024).
@@
where $v_{\perp}$ corresponds to the relative velocity of the source-lens in the direction perpendicular to the line-of-sight of the observer (and in our case is related to the orbital velocity of the star). The presence of a DH would affect $\theta_{E}$, and possibly the velocity $v_{\perp}$ if the trajectory of the source is modified due to the difference in the gravitational potential and/or the existence of dynamical friction effects. We will not explore here such effects, and leave them for future research. It is important to notice that the physical quantities which enter into the Einstein time cannot be measured by lensing alone, and more information is needed. Astrometry observations can indeed break the degeneracy between these parameters through a measurement of the velocity. To model the actual light curve associated with the lensing effect in the time-domain we assume a trajectory for the source as
@@
Assuming a Keplerian orbit for the light source around the DH lens (i.e., $v_{\perp}\approx\sqrt{\frac{GM}{A}}$, with $A$ the semi-major axis of the orbit which is $A\approx D_{\rm LS}$) and under our main assumption that $D_{\rm L}\approx D_{\rm S}\gg D_{\rm LS}$, the Einstein time is $t_{E}=2D_{\rm LS}/c$. Using Kepler’s laws the Einstein time can be expressed as a function of the period $T$ and the total mass of the system $M$ as
@@
For a system like the Gaia BH1, the total mass is $\sim 11M_{\odot}$, the mass ratio between the total mass and the luminous-star mass is $\sim 14$, the period is $185.6$ d, and the size of the orbit is $\sim 1.4AU$ (El-Badry et al., 2022). For such a system the Einstein time is of the order of $\sim 23$ min. The time evolution of the apparent magnitude of the source as it passes behind the lens can be seen in Figure 4 for the case of two occultation events, the first with $u_{0}=0$ and the second with $u_{0}=0.91x_{R}$, respectively. Although the orientation of the BH1 system is not favorable for such an observation, other similar systems may be optimally oriented.
Figure 4: Similar $\Delta m$ to Figure 3, but now as a function of time. The evolution depends on the characteristic time $t_{E}$ (see (13)), and we have assumed the same characteristics as those of the Gaia BH1 system. The figure shows two transit events, one with initial impact parameters $u_{0}=0$ and $u_{0}=0.91x_{R}$, respectively. The transit for the first event lasts a total of $\sim 40$min, while the spikes in the magnification that happen as the source passes behind the edge of the DH lens last $\sim 5$min each. For the rest of the transit, the source is brighter than its normal brightness and behaves like a point mass. The second transit lasts less time, $\sim 25$min since the light source grazes the lens. In this case, we observe two magnification spikes as the source goes behind the lens, while the source remains magnified throughout the transit. This plot should be compared with similar ones for e.g. by Romao and Croon (2024).
IV Dark halo light curve (Black hole with a dark halo)